# Direct measurement of atomic number density using Single Pass Absorption Spectroscopy (SPAS)

### v1

@@ II.4 Light propagation through the vapor cell
The propagation of resonant light through an atomic vapor is governed by absorption and dispersion processes that modify the transmitted intensity along the direction of propagation. For a monochromatic beam incident on a homogeneous vapor of length $z$, the transmitted intensity follows the Beer Lambert law,
@@
where $I_{0}$ is the incident intensity, and $\alpha^{\prime}_{ij}(\Delta,T)$ is the Doppler broadened absorption coefficient (from Equation 9), which depends on the detuning $\Delta$ and the vapor temperature $T$. The absorption coefficient encapsulates the combined effects of the atomic line profile including natural, transit time, and Doppler broadening.
The normalized transmission through the vapor cell is thus expressed as
@@
The simulated transmission spectrum obtained from Equation 11 exhibits good agreement with experimental data, and nonlinear least-squares fitting is employed to extract physical parameters atomic number density.
@@ III Experimental setup
We performed single-pass absorption spectroscopy [14, 11] on naturally abundant rubidium vapor cells (72.2% 85Rb, 27.8% 87Rb; 100 mm length) and MEMS-based vapor cells (2 mm length) to study their absorption characteristics at the $D_{2}$ transition (780.24 nm), as shown in Figure 1. A schematic of the experimental setup is presented in Figure 2.
Figure 2: Schematic of the experimental setup for absolute absorption spectroscopy and frequency calibration. The absorption signal transmitted through the rubidium vapor cell is detected using photodetector (PD-1), while the reference transmission signal from the Fabry–Perot interferometer (FPI) is monitored by photodetector (PD-2). Both signals are simultaneously recorded with a digital storage oscilloscope (DSO). The piezoelectric transducer (PZT) scan signal from the external cavity diode laser (ECDL) controller is used as the trigger input to synchronize the oscilloscope trace with the frequency scan. Optical components include mirrors (M1–M5), neutral density filters (NDF), lenses (L1–L3), polarizing beam splitters (PBS), an electro-optic modulator (EOM), and a beam dump (BD). The vapor cell temperature is stabilized using a temperature controller (TEC) and resistive foil heaters monitored by four NTC (Negative Temperature Coefficient) temperature sensors.
A tunable external cavity diode laser (Toptica DL PRO 780) operating at $780.24$ nm, equipped with an inbuilt $35$ dB optical isolator, was used as the light source. Two three-axis mirrors (M1 and M2) are used to align the laser light horizontally and vertically. The laser beam was initially split into two paths using a half-wave plate (HWP) and a polarizing beam splitter (PBS). One beam was directed toward a frequency reference measurement, while the other was used for absorption spectroscopy. The intensity of the spectroscopy beam was further adjusted using an additional combination of HWP and PBS.
@@
The vapor cell was enclosed in a custom-fabricated heating assembly integrated with a temperature controller (Thorlabs TC300B) to facilitate spectroscopic measurements over a range of temperatures. However, the temperature displayed by the controller does not accurately reflect the actual temperature of the vapor cell due to thermal gradients in the heating assembly. Four precalibrated temperature sensors were attached at different locations in the cell to accurately determine the cell temperature. Two sensors were placed at the opposite edges and two at the center of the vapor cell (as illustrated in Figure 2). The temperature readings from these sensors were recorded using an Arduino Uno, and the average value was considered as the effective vapor cell temperature for subsequent simulations. The absorption signal was detected using an unbiased photodetector (Thorlabs PDA36A2) and recorded on one channel of a digital storage oscilloscope (Tektronix MSO44, 4-channel). On the other hand, for the frequency reference measurement, the second beam was passed through an electro-optic modulator (Qubig PM7_\_NIR_\_25) driven by a $25$ MHz radiofrequency (RF) signal, which modulated the frequency of the laser beam. The modulated beam was then collimated using a lens system (L1 and L2) and directed through a mode-matching lens (L3) before entering a Fabry-Perot interferometer (Thorlabs SA30-73 with FSR=$1500$ MHz). The transmitted signal from the interferometer, which included sidebands resulting from the 25 MHz modulation, was captured by another photodetector and recorded on a separate channel of the oscilloscope. This signal provides a precise reference for frequency calibration. The rubidium absorption signal and the interferometer transmission peaks were acquired simultaneously within the same oscilloscope time window, ensuring synchronized data acquisition for accurate analysis.
@@
For quantitative analysis of the absolute absorption measurement from the measured transmission spectra, it is essential to determine the offset level of the photodetector. This offset mainly arises from the intrinsic dark current of the photodetector. Before measuring the transmission data, we measured the photo detector offset voltage by keeping the probe light completely off and closing the input port so that no stray light could fall on the photo detector. This offset voltage is used as the zero level where no light falls on the photodetector for all the subsequent absorption measurements. Another way to measure the zero level is to optically opaque the atomic vapor by significantly increasing its temperature so that no probe light is transmitted through the medium [31]. Although this technique offers a physical reference point for complete absorption, it is not universally applicable. In particular, for miniaturized vapor cells (MEMS cells) with inherently low atomic density or for investigating any weak transitions that require extreme heating to achieve complete opacity, this technique becomes difficult to implement. In such scenarios, the dark-current-based method offers a more flexible and non-invasive strategy for baseline normalization. Instead of this baseline correction, two other nonlinearities are present in the measured absorption spectra due to the piezoelectric actuator. A piezoelectric actuator is used to tune the laser frequency, which introduces inherent nonlinearities in the laser power and frequency of the measured transmission spectra. To remove these non-linear effects from the frequency and amplitude of the measured transmission spectrum, we adopted the methodology as prescribed in [31]. The frequency non-linearity is removed using the Fabry-Perot interferometer (FPI) signal, and the power nonlinearity is removed using the off-resonant region of the measured transmitted spectrum.
@@ IV Results and discussion
In this study, we systematically measure absorption spectroscopy of rubidium vapor under various experimental conditions to determine the effect of probe power, cell temperature and cell length. All of these parameters are precisely measured and incorporated into the theoretical model, and the model spectra were then fitted to the experimental data by treating the number density as a free parameter to achieve quantitative agreement.
For a $100$ mm long vapor cell containing naturally abundant rubidium (72.2% 85Rb and 27.8% 87Rb), single-pass absorption spectra were recorded at three different temperatures: $20.3\pm 0.3^{\circ}$C, $39.5\pm 0.8^{\circ}$C, and $66.5\pm 3.3^{\circ}$C. The corresponding experimental and theoretically fitted spectra, along with their residuals, are shown in Figure 3. Across all temperature datasets, the fits exhibited an $R^{2}$ (coefficient of determination) value that exceeded 0.99, confirming the accuracy of the theoretical model in reproducing the experimental line shapes.
Figure 3: Experimental and fitted absorption spectra as a function of detuning for three different temperatures: $20.3^{\circ}$C (red solid line for fit, blue dots for experiment), $39.5^{\circ}$C (black dashed line for fit, orange dots for experiment), and $66.5^{\circ}$C (blue dotted line for fit, green dots for experiment). The lower panels show the corresponding residuals (experimental minus fitted). The fitted spectra exhibit excellent agreement with experimental data across all temperatures, with residuals remaining within $\pm 0.25$ in absolute transmission.
To further validate the generality of the model, absorption spectroscopy was also performed on isotopically enriched rubidium vapor cells: an $86.5\%$ enriched 85Rb cell and a $93.6\%$ enriched 87Rb cell, each with a length of $100$ mm. Spectra were recorded at the same three cell temperatures ($20.3\pm 0.3^{\circ}$C, $39.5\pm 0.8^{\circ}$C, and $66.5\pm 3.3^{\circ}$C). For both isotopic compositions, the experimental spectra were well reproduced by the theoretical model, yielding $R^{2}>0.99$ in all cases. This consistency highlights the versatility and reliability of the model across different isotopic mixtures.
Furthermore, to verify the validation of our model beyond the weak-probe limit [32], we perform the absorption spectroscopy measurement at a higher probe intensity ($\sim 2~I_{\text{sat}}$). Appendix Figure 8 illustrates the very good agreement ($R^{2}>0.99$) between the experimentally measured and theoretically fitted spectra from our model (see Appendix Appendix C for more details).
The model was further tested using a micro-electro-mechanical systems (MEMS) vapor cell with an optical path length of 2 mm, representing compact vapor cells relevant for chip-scale atomic devices and portable quantum sensors. Figure 4 presents the comparison between simulated and experimental transmission spectra at three operating temperatures: $34.2\pm 0.3^{\circ}$C, $51.1\pm 0.8^{\circ}$C, and $59.1\pm 1^{\circ}$C. The fitted spectra showed excellent agreement with the experimental data, with residuals remaining below $0.05$ in absolute transmission, indicating that the model successfully captures the dominant physical mechanisms even in strongly confined geometries.
Figure 4: Comparison of simulated and experimental transmission spectra of the rubidium MEMS vapor cell at $34.2^{\circ}$C, $51.1^{\circ}$C, and $59.1^{\circ}$C. Lower panels show the residuals, demonstrating excellent agreement across the temperature range.
@@
The atomic number density of rubidium vapor can be estimated theoretically from its equilibrium vapor pressure and experimentally from absorption spectroscopy measurements. The temperature dependence of the vapor pressure of alkali metals has been comprehensively studied, with reliable empirical relations reported by Alcock et al. [25] over a wide temperature range (298–2500 K). For rubidium, the vapor pressure $P_{\mathrm{vap}}(T)$ (in torr) as a function of temperature $T$ (in Kelvin) is given by
$\log_{10}P_{\mathrm{vap}}(T)=2.881+a-\frac{b}{T},$ (12)
where the constants $a$ and $b$ correspond to different physical phases: $a=4.857,b=4215~\text{(solid phase)},a=4.312,b=4040~\text{(liquid phase)}.$ Here, $P_{\mathrm{vap}}$ is expressed in torr; when expressed in atmospheres, the additive constant 2.881 is omitted. The corresponding atomic number density $N(T)$ is determined from the ideal gas relation,
$N(T)=\frac{P_{\mathrm{vap}}(T)}{k_{\mathrm{B}}T},$ (13)
where $k_{\mathrm{B}}$ is the Boltzmann constant and $P_{\mathrm{vap}}$ is converted to Pascals using $1~\text{torr}=133.322~\text{Pa}$. This formulation, first proposed by Killian [33], remains a standard approach for estimating alkali vapor densities and provides a useful benchmark for experimental comparison.
From the experimental absorption spectra, the atomic number density was extracted by fitting the data using our theoretical model. To verify the accuracy of the measured number density, the obtained values were compared with those estimated from the empirical relation Equation 12 provided by Alcock et al. for rubidium vapor pressure and number density as a function of temperature (Equation 13). The results for both the $100$ mm vapor cell and the $2$ mm MEMS cell are presented in Figure 5. The experimentally determined densities show excellent agreement with the empirical vapor pressure model of Alcock et al. [25] across the investigated temperature range of 293293343343 K. This agreement validates the reliability of our absorption-based method for determining the absolute number density of rubidium vapor under varying thermal and geometrical conditions.
Figure 5: Rubidium vapor number density as a function of temperature obtained from the 100 mm cell (purple circles) and MEMS cell (red diamonds), compared with the empirical model of Alcock et al.  [25] (blue dashed line). The vertical red dashed line marks the solid–liquid phase transition at 312.45 K. Insets (a)–(c) highlight the temperature ranges 291–295 K, 311–315 K, and 339–342 K, respectively, showing excellent agreement between the experimentally measured and theoretically predicted number densities within experimental uncertainties.
Although the NTC temperature sensors used in our experiment setup have an intrinsic accuracy of $\pm 0.3^{\circ}$C, the observed error bars in the number density of the 100 mm long vapor cell at higher temperatures are slightly larger than at lower temperatures. At higher temperatures, we observe that the rubidium atoms inside the vapor cell tend to deposit on the inner walls through which the laser beam is passing. This introduces abrupt non-linearity in the absorption signal. To mitigate this issue, we introduced a slight temperature gradient along the cell in such a way that the flat glass window through which the laser beam passes remains slightly hotter than the center cylindrical region. Although this configuration helps us suppress the condensation on the flat window, it marginaly increases the uncertainty in the temperature measurement.
@@
A more accurate temperature control was achieved for the MEMS vapor cell, where the cell was mounted within a machined aluminum block equipped with four sensors placed symmetrically around the cell edges. This configuration ensured uniform heating and minimized temperature gradients. As shown in Figure 5, the error bars in the MEMS cell measurements are substantially smaller, indicating improved temperature uniformity and more precise determination of the number density. These observations emphasize that precise temperature monitoring and control play a crucial role in reducing uncertainty in vapor density measurements, particularly at higher operating temperatures.
@@
The close correspondence between the experimental data and the theoretical model across different cell geometries underscores the consistency and robustness of our approach. In particular, the successful reproduction of the empirical trend using both standard ($100$ mm) and MEMS ($2$ mm) vapor cells demonstrates the scalability of the absorption-based number density determination method. This result is significant for precision spectroscopy and the development of compact, chip-scale atomic devices, where accurate control of vapor density is crucial for optimal performance.
@@ V Conclusion
We have developed a theoretical model that accurately predicts the atomic number density of the Rb vapor cells with different isotopic compositions and cell lengths, if the experimental parameters such as the temperature of the vapor cell (at different points), laser beam diameter, and optical power are provided. The model effectively shows the dependence of absorption spectra on both laser power and cell length.
Determining the absolute transmission accurately from the experiment is essential for achieving quantitative agreement between theory and experiment. In this work, we have carefully accounted for the photodetector’s offset voltage (which primarily arises from its intrinsic dark current). The dark current approach provides a reliable and noninvasive baseline correction, particularly valuable when complete absorption cannot be achieved even at elevated temperatures. We used it as the zero-reference level for all absorption data. The alternative methods, such as heating the vapor to increase the opacity of the medium, offer the physical reference points. Nevertheless, they are not always practical, such as in miniaturized vapor cells or when studying weak transitions. The dark current approach provides a reliable and non invasive baseline correction, particularly valuable when complete absorption cannot be achieved. The theoretical model described in this work can be extended to predict absorption spectra and to calculate the atomic number density of other elements or molecules, provided that the energy level diagrams are known. Future work could also include additional transitions in the Rb atom to establish a complete formalism for the saturation absorption spectroscopy. These extensions will enhance the model’s versatility in atomic physics applications and the development of quantum technology.
@@
Sumit Achar and Shivam Sinha contributed equally to this work as co-first authors, theoretical model development, experimental design, data analysis and manuscript writing. Ezhilarasan M. and Chandankumar R.: experimentation, data collection and analysis support. Arijit Sharma: conceptualization of the project, manuscript writing, review, editing, and overall project supervision.

### v2

@@ II.4 Light propagation through the vapor cell
The propagation of resonant light through an atomic vapor is governed by absorption and dispersion processes that modify the transmitted intensity along the direction of propagation. For a monochromatic beam incident on a homogeneous vapor of length $z$, the transmitted intensity follows the Beer-Lambert law,
@@
where $I_{0}$ is the incident intensity, and $\alpha^{\prime}_{ij}(\Delta,T)$ is the Doppler-broadened absorption coefficient (from Equation 11), which depends on the detuning $\Delta$ of laser and the vapor temperature $T$. The absorption coefficient includes the combined effects of the atomic line profile, natural, transit time, and Doppler broadening.
Spectroscopic measurements are performed by monitoring the fraction of light transmitted through the vapor cell, and transmission provides a dimensionless and experimentally observable quantity that reflects the frequency-dependent absorption of the medium. It is therefore convenient to express the light–matter interaction in terms of the normalized transmission. The normalized transmission through the vapor cell is expressed as,
@@ III Experimental setup
We performed single-pass absorption spectroscopy [25, 26] on naturally abundant rubidium vapor contained in both a conventional vapor cell of length 100 mm and a microfabricated MEMS-based vapor cell of length 2 mm. The measurements were carried out primarily on the $D_{2}$ transition of rubidium at 780.24 nm, as shown in Figure 2. A schematic of the experimental setup is presented in Figure 3.
Figure 3: Schematic of the experimental setup for absolute absorption spectroscopy and frequency calibration. The absorption signal transmitted through the rubidium vapor cell is detected using photodetector (PD-1), while the reference transmission signal from the Fabry–Pérot interferometer (FPI) is monitored by photodetector (PD-2). Both signals are simultaneously recorded with a digital storage oscilloscope (DSO). The piezoelectric transducer (PZT) scan signal from the external cavity diode laser (ECDL) controller is used as the trigger input to synchronize the oscilloscope trace with the frequency scan. Optical components include mirrors (M1–M5), neutral density filters (NDF), lenses (L1–L3), polarizing beam splitters (PBS), an electro-optic modulator (EOM), and a beam dump (BD). The vapor cell temperature is stabilized using a temperature controller and resistive foil heaters, which are monitored by four NTC (Negative Temperature Coefficient) temperature sensors.
A tunable external-cavity diode laser (ECDL, Toptica DL PRO 780) operating at $780.24$ nm, equipped with an inbuilt $35$ dB optical isolator, was used as the light source. The output beam was spatially aligned using three-axis mirrors (M1 and M2) and subsequently divided into a spectroscopy arm and a frequency-reference arm using a half-wave plate (HWP) and a polarizing beam splitter (PBS). The optical power in the spectroscopy arm was adjusted using an additional HWP-PBS combination and a neutral density filter (NDF).
@@
The vapor cell was enclosed in a custom-fabricated heating assembly integrated with a temperature controller (Thorlabs TC300B) to facilitate spectroscopic measurements over a range of temperatures. However, the temperature displayed by the controller does not accurately reflect the actual temperature of the vapor cell due to thermal gradients in the heating assembly. Four precalibrated NTC temperature sensors were attached at different locations in the cell to accurately determine the cell temperature (as illustrated in Figure 3). The temperature readings from these sensors were recorded using an Arduino Uno, and the average value was considered as the effective vapor cell temperature for subsequent simulations. The absorption signal was detected with an unbiased photodetector (PD-1, Thorlabs PDA36A2) and recorded on one channel of a 4-channel digital storage oscilloscope (Tektronix MSO44B).
On the other hand, for the frequency-reference measurement, the second beam was passed through an electro-optic modulator (Qubig PM7_\_NIR_\_25) driven by a $25$ MHz radio-frequency (RF) signal, which modulated the frequency of the laser beam. The modulated beam was then collimated using a lens system (L1 and L2) and directed through a mode-matching lens (L3) before entering a Fabry-Pérot interferometer (Thorlabs SA30-73 with FSR= $1500$ MHz). The transmitted signal from the interferometer, including sidebands from the 25 MHz modulation, was captured by another photodetector (PD-2) and recorded on a separate channel of the oscilloscope. This signal provides a precise reference for frequency calibration. The rubidium absorption signal and the interferometer transmission peaks were acquired simultaneously within the same oscilloscope time window, ensuring synchronized data acquisition for accurate analysis.
@@
For quantitative analysis of absolute absorption, it is essential to calibrate the baseline of the detected transmission signal. The baseline offset arises from the intrinsic dark current of the photodetector. This offset voltage was measured independently by blocking the probe beam and optically isolating the detector from ambient light. The measured offset was subsequently subtracted from all recorded transmission signals and used to define the zero-transmission reference level. Another way to measure the zero level is to make the atomic vapor optically opaque by significantly increasing its temperature, so that probe light is not transmitted through the medium [45]. Although this technique offers a physical reference point for complete absorption, it is not universally applicable. In particular, for miniaturized vapor cells (MEMS cells) with inherently low atomic density or for investigating any weak transitions that require extreme heating to achieve complete opacity, this technique becomes difficult to implement. In such scenarios, the dark-current-based method offers a more flexible and non-invasive strategy for baseline normalization.
Non-linearities in both the laser frequency scan and the output power, introduced by the piezoelectric actuator used for frequency tuning, were corrected by the methodology prescribed in [45]. The frequency non-linearity was removed by mapping the Fabry-Pérot interferometer (FPI) transmission peaks to their known frequency separations, and the power non-linearity was removed using the off-resonant region of the measured transmitted spectrum. The off-resonant regions were fit with a higher-order polynomial to characterize the power variation across the scan, which was subsequently used to normalize the measured absorption signal.
In addition to the absolute absorption measurement at 780.24 nm, absorption spectroscopy was also performed on the $5S_{1/2}\to 6P_{3/2}$ transition at 420.29 nm using the same experimental methodology as shown in Figure 3. The optical layout and measurement protocol remained unchanged, with wavelength-appropriate substitutions for the laser source, photodetectors, and coated optical components. All data acquisition, baseline correction, and frequency calibration procedures were identical to those employed at 780.24 nm. It enables a direct comparison of number-density estimates obtained from various transitions with different strengths and saturation intensities.
@@ IV Results and discussion
We systematically measured the single-pass absorption spectrum of rubidium vapor under various experimental conditions to determine the effects of probe power, cell temperature, and cell length. All of these parameters are precisely measured and incorporated into the theoretical model, and the model spectra were then fitted to the experimental data by treating the number density as a free parameter to achieve quantitative agreement.
@@
The atomic number density of the rubidium vapor can be estimated theoretically from its equilibrium vapor pressure. The temperature dependence of the vapor pressure of alkali metals has been comprehensively studied, with reliable empirical relations reported by Alcock et al. [20] over a wide temperature range (298–550 K). From the experimental absorption spectra, we extracted the atomic number density by fitting the data with our theoretical model. To verify the accuracy of the measured number density, the obtained values were compared with those estimated from the empirical relation Equation 1 provided by Alcock et al. for rubidium vapor pressure and number density as a function of temperature (see Equation 2). This comparison provides a stringent consistency check of both the experimental procedure and the theoretical modeling.
Figure 4: Comparison of simulated and experimental transmission spectra of the rubidium MEMS vapor cell at $307.35$ K, $324.25$ K, and $332.25$ K. Lower panels show the residuals (experiment-simulation).
@@
For the MEMS vapor cell (length 22~$m$) containing naturally abundant rubidium (72.2% 85Rb and 27.8% 87Rb), single-pass absorption spectra were recorded at various temperatures: ranging from 293363293-363K using a probe intensity of approximately $\sim 0.2~I_{\text{sat}}$. The corresponding experimental and theoretically fitted spectra, along with their residuals, are shown in Figure 6. The fitted spectra yield coefficients of determination $R^{2}>0.99$ for all temperatures, indicating excellent quantitative agreement. The cell temperature was monitored using precalibrated NTC temperature sensors with an intrinsic accuracy of $\pm~0.3$ K. The cell was mounted in a machined aluminum block equipped with four symmetrically placed sensors to ensure uniform heating. At elevated temperatures, a residual temperature gradient of approximately 1–2 K was observed, which accounts for the slightly larger error bars in the extracted number density at higher temperatures. As shown in Figure 6, the error bars in the MEMS cell measurements are substantially smaller, indicating the improved temperature uniformity and more precise determination of the number density. These observations emphasize that precise temperature monitoring and control are crucial for reducing uncertainty in vapor density measurements, particularly at higher operating temperatures.
We performed the same set of measurements with a commercial 100 mm Rb vapor cell with both wavelengths, 780 nm and 420 nm. Since the atomic number density is a thermodynamic property of the vapor, it is independent of the optical interaction length. The results presented in Appendix Appendix A, confirm that the same density values are obtained for both the MEMS and the standard-length cells.
Furthermore, to verify the validation of our model beyond the weak-probe regime [46], we perform the absorption spectroscopy measurement at a higher probe intensity ($\sim 2~I_{\text{sat}}$). As shown in Appendix Appendix B (see Figure 11 for more details), the model continues to reproduce the experimental spectra with $R^{2}>0.99$, confirming its robustness over a wider range of excitation conditions.
In addition to spectroscopy on the $D_{2}$ transition ($5S_{1/2}\rightarrow 5P_{3/2}$) at 780.24 nm, we also performed measurements using a $420~\mathrm{nm}$ laser addressing the $5S_{1/2}\rightarrow 6P_{3/2}$ transition. The number density is an intrinsic thermodynamic property of the rubidium vapor, determined solely by the vapor pressure at a given temperature, and is independent of the specific excited state used for optical interrogation. The agreement in the extracted number density values from both transitions thus provides a strong internal consistency check. For a given vapor cell temperature, the extracted atomic number density remains identical irrespective of whether the atoms are excited via the $5S_{1/2}\rightarrow 5P_{3/2}$ transition or the $5S_{1/2}\rightarrow 6P_{3/2}$ transition.
Figure 5: Experimental and simulated transmission spectra comparing the 780 nm and 420 nm rubidium transitions at the same optical power of $0.2\,~I_{\mathrm{sat}}$ and nearly identical temperatures (333.25 K and 332.25 K). The upper panel shows simulated spectra (solid red: 780 nm, dashed black: 420 nm) with corresponding experimental data (markers). The lower panels present the residuals.
Figure 5 compares representative transmission spectra recorded at 780 nm and 420 nm under nearly identical temperature conditions and at the same normalized probe intensity ($0.2,I_{\mathrm{sat}}$). As expected, the absorption strength of the 420 nm transition is significantly weaker than that of the $D_{2}$ transitions. The successful application of the model to MEMS cells has important implications for the development of chip-scale atomic devices, where precise knowledge of the atomic number density is crucial for optimizing sensor performance. The ability to accurately determine the vapor density in these miniaturized systems enables more accurate calibration and characterization of MEMS-based atomic, quantum, and precision timing sensors.
The extracted number density of rubidium in the MEMS cell as a function of temperature is shown in Figure 6, where results obtained from 780 nm and 420 nm absorption measurements over a range of temperatures 293363293-363 K are compared with the empirical relation from Equation 1. The measured number densities follow the empirical vapor-pressure relation closely over the entire temperature range, with deviations well within the experimental uncertainties. Although absorption data were acquired over a broad temperature range at both probe wavelengths, only representative data points at selected temperatures are displayed in the figure to maintain visual clarity.
Figure 6: Rubidium vapor number density as a function of temperature for the MEMS vapor cell. Experimental number densities extracted from absorption measurements at 780 nm (red diamonds) and fluorescence detection at 420 nm (blue diamonds) are compared with the empirical vapor pressure model (dashed line). The vertical dashed line indicates the phase transition temperature at 312.45 K. Insets (a), (b), and (c) show magnified views around 298 K, 313 K, and 332 K, respectively.
@@
The successful reproduction of the empirical vapor-pressure curve using both standard ($100$ mm) and MEMS ($2$ mm) vapor cells demonstrates the scalability of the absorption-based number density determination method. This result is significant for precision spectroscopy and the development of compact, chip-scale atomic devices, where accurate control of vapor density is crucial for optimal performance.
@@ V Conclusion
We have presented a quantitative method for extracting the absolute atomic number density of warm rubidium vapor in chip-scale MEMS vapor cells using single-pass absorption spectroscopy. The approach is based on a multi-level Lindblad master equation formalism that incorporates hyperfine structure, Doppler averaging, optical pumping, and transit-time relaxation. All experimentally measured parameters, including vapor-cell temperature, optical path length, beam diameter, and optical power, are directly included in the model.
The method was experimentally validated using a MEMS vapor cell over a broad temperature range. Despite the intrinsically low optical depth associated with such short path lengths, the extracted number densities exhibit agreement with the empirical vapor-pressure relation of Alcock et al.. Consistent density values were obtained from absorption measurements on both the strong $5S_{1/2}\!\rightarrow\!5P_{3/2}$ (780.24 nm) transition and the weak $5S_{1/2}\!\rightarrow\!6P_{3/2}$ (420.29 nm) transition shows the applicability of the technique in weak-absorption regimes relevant to MEMS-scale devices.
The intrinsic photodetector dark current was used as a zero-transmission reference to provide a stable, noninvasive calibration method well suited to MEMS vapor cells, where complete absorption cannot be achieved without impractically high temperatures. The resulting fits reproduce the measured spectra with high fidelity ($R^{2}>0.99$). The demonstrated framework enables precise and reproducible characterization of vapor density in MEMS alkali cells.
The theoretical model described in this work may be extended to predict absorption spectra and to calculate the atomic number density of other atomic or molecular vapors, provided that the energy-level diagrams are known. Future work could also include additional transitions in the Rb atom to establish a complete formalism for the saturation absorption spectroscopy. These extensions will enhance the model’s versatility in atomic physics applications and the development of quantum technology.
@@
Sumit Achar and Shivam Sinha contributed equally to this work as co-first authors, theoretical model development, experimental design, data analysis, and manuscript writing. Ezhilarasan M. and Chandankumar R.: experimentation, data collection and analysis support. Arijit Sharma: conceptualization of the project, manuscript writing, review, editing, and overall project supervision.